slipknot3066:

Write the equation of a circle with a center at (1, 4) where a point on the circle is (4, 8).

slipknot3066:

@Narad

Narad:

The centre of the circle is (1,4) The radius of the circle is \[=\sqrt{((4-1)^2+(8-4)^2)}\] \[=\sqrt{(9+16)}=\sqrt{(25)}=5\] The equation of the circle is \[(x-h)^2+(y-k)^2=r^2\] \[(x-1)^2+(y-4)^2=25\]